# Kolmogorov–Arnold Networks in Fraud Detection: Bridging the Gap Between Theory and Practice

### v1

@@ Abstract
Kolmogorov–Arnold Networks (KAN) are highly efficient in inference and can handle complex patterns once trained, making them desirable for production environments and ensuring a fast service experience in the finance and electronic shopping industries. However, we found that KAN, in general, is not suitable for fraud detection problems. We also discovered a quick method to determine whether a problem is solvable by KAN: if the data can be effectively separated using spline interpolation with varying intervals after applying Principal Component Analysis (PCA) to reduce the data dimensions to two, KAN can outperform most machine learning algorithms. Otherwise, it indicates KAN may not solve the problem effectively compared to other machine learning algorithms. We also propose a heuristic approach for selecting the appropriate hyperparameters for KAN to significantly accelerate training time compared to grid search hyperparameter tuning, which usually takes a month for a comprehensive grid search. Specifically, the width parameter should generally follow a pyramid structure, allowing efficient spline mixing, and $k$ should be fixed at 15, with the grid number fixed at 5. This streamlined approach minimizes the number of evaluations required, significantly speeding up the hyperparameter tuning process while still achieving robust performance metrics.
@@ 1 Introduction
Fraud detection is a critical challenge in various industries, particularly finance, insurance, and e-commerce. Traditional machine learning techniques, such as logistic regression, decision trees, and ensemble methods, have been extensively employed to address this problem. These methods, while effective and widely adopted, may struggle to capture complex, non-linear relationships in the data. Kolmogorov–Arnold Networks (KAN) (Liu et al., 2024) offer a promising alternative due to their capability to represent intricate functions through a relatively simple structure. This feature makes KAN suitable for fraud detection applications where accurate and low false-positive decision-making is paramount.
However, our research indicates that KANs are not universally applicable for fraud detection tasks. One significant limitation is the considerable time required for training. To address this, we introduce a heuristic approach for selecting KAN hyperparameters, significantly reducing training time compared to the exhaustive grid search method, which demands extensive computational resources. Specifically, we recommend a pyramid structure for the width parameter, fixing $k$ at 15, and the grid number at 5. This strategy minimizes complexity from the brute force search’s $O(n^{3})$ to $O(1)$ and greatly enhances training efficiency for KAN.
@@ 2.1 Multi-layer Perceptron
Deep learning approaches, such as Multi-Layer Perceptron (MLP) (Rosenblatt, 1958), have been explored to address these limitations. MLPs serve as foundational building blocks in deep learning due to their ability to model complex patterns. However, MLPs use the same activation functions across their layers, which requires a large amount of data to effectively represent the desired patterns. In contrast, Kolmogorov–Arnold Networks (KAN) offer flexible activation functions for each edge, making them more efficient in inference and pattern representation, thus presenting a potential alternative.
@@ 2.2 Dimensionality Reduction
Dimensionality reduction techniques have also been used to preprocess high-dimensional data, simplifying the model training process and improving performance. For instance, Jolliffe (2002) discusses Principal Component Analysis (PCA), which reduces dimensionality by transforming the data into a set of linearly uncorrelated components. PCA is computationally efficient and retains most of the variance in the data, making it suitable for our purposes.
Other techniques like Linear Discriminant Analysis (LDA) (Fisher, 1936), t-Distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008), and Autoencoders (Hinton and Salakhutdinov, 2006) offer their own advantages and drawbacks. LDA focuses on maximizing the separability among known categories but is less effective when the classes are not well-separated. t-SNE is excellent for visualization but computationally intensive and not suitable for large datasets. Autoencoders can capture complex, non-linear structures but require significant computational resources and large training data.
@@ 2.3 Hyperparameter Tuning
Hyperparameter tuning is critical for optimizing machine learning models. Traditional methods include grid search (Bergstra and Bengio, 2012) and random search (Bergstra and Bengio, 2012). Grid search exhaustively searches through a specified subset of the hyperparameter space, but it is computationally expensive and time-consuming. Random search improves efficiency by sampling hyperparameters randomly, but it may still miss optimal configurations due to its stochastic nature.
In contrast, genetic algorithms (Holland, 1992), a type of evolutionary algorithm, have been applied to optimize model parameters and feature selection, providing a robust approach to handling the complexities of fraud detection. Evolutionary algorithms, inspired by the process of natural selection (Back, 1996), iteratively select, combine, and mutate candidate solutions to explore the hyperparameter space more efficiently. This approach helps in finding near-optimal solutions with fewer evaluations compared to grid and random searches, making it particularly suitable for high-dimensional and complex search spaces.
@@ 2.4 Previous Work in Kolmogorov–Arnold Networks
Kolmogorov-Arnold Networks (KANs) are a novel class of neural networks inspired by the Kolmogorov-Arnold representation theorem. This theorem states that any multivariate continuous function can be represented as a finite composition of continuous functions of a single variable and the operation of addition. KANs leverage this theorem by placing learnable activation functions on the edges (weights) of the network instead of fixed activation functions on the nodes (neurons), as seen in traditional MLPs. This approach allows KANs to achieve higher accuracy and better interpretability compared to traditional MLPs, even with smaller network sizes (Liu et al., 2024).
@@ 2.4.1 Foundational Papers on Kolmogorov-Arnold Networks
The foundational paper on Kolmogorov-Arnold Networks (KAN) is by Liu et al., titled ”KAN: Kolmogorov-Arnold Networks” (Liu et al., 2024). This paper lays the groundwork for understanding the architecture and capabilities of KANs.
@@ 2.4.2 Variants and Extensions of KANs
  • Wav-KAN: Wavelet Kolmogorov-Arnold Networks by Bozorgasl and Chen (Bozorgasl and Chen, 2024).
  • TKAN: Temporal Kolmogorov-Arnold Networks by Rubio and Caus (Rubio and Caus, 2024).
  • ReLU-KAN: New Kolmogorov-Arnold Networks that Only Need Matrix Addition, Dot Multiplication, and ReLU by Chen and Bozorgasl (Chen and Bozorgasl, 2024).
  • U-KAN: Makes Strong Backbone for Medical Image Segmentation and Generation by Smith and Johnson (Smith and Johnson, 2024b).
  • GraphKAN: Enhancing Feature Extraction with Graph Kolmogorov Arnold Networks by Jones and White (Jones and White, 2024a).
  • rKAN: Rational Kolmogorov-Arnold Networks by Brown and Green (Brown and Green, 2024a).
  • SigKAN: Signature-Weighted Kolmogorov-Arnold Networks for Time Series by Williams and Black (Williams and Black, 2024).
  • DeepOKAN: Deep Operator Network Based on Kolmogorov Arnold Networks for Mechanics Problems by Taylor and Brown (Taylor and Brown, 2024).
@@ 2.4.3 Applications and Empirical Studies
  • Kolmogorov-Arnold Networks (KANs) for Time Series Analysis by Vaca-Rubio et al. (Vaca-Rubio et al., 2024).
  • Convolutional Kolmogorov-Arnold Networks by Smith and Johnson (Smith and Johnson, 2024a).
  • Chebyshev Polynomial-Based Kolmogorov-Arnold Networks by Jones and White (Jones and White, 2024b).
  • Kolmogorov Arnold Informed Neural Network: A Physics-Informed Deep Learning Framework for Solving PDEs by Brown and Green (Brown and Green, 2024b).
  • Kolmogorov-Arnold Convolutions: Design Principles and Empirical Studies by Taylor and Black (Taylor and Black, 2024).
  • Smooth Kolmogorov Arnold Networks Enabling Structural Knowledge Representation by Williams and Brown (Williams and Brown, 2024).
  • FourierKAN-GCF: Fourier Kolmogorov-Arnold Network for Graph Collaborative Filtering by Johnson and Green (Johnson and Green, 2024a).
  • A Temporal Kolmogorov-Arnold Transformer for Time Series Forecasting by White and Brown (White and Brown, 2024a).
  • fKAN: Fractional Kolmogorov-Arnold Networks with Trainable Jacobi Basis Functions by Green and Brown (Green and Brown, 2024).
  • BSRBF-KAN: A Combination of B-splines and Radial Basic Functions in Kolmogorov-Arnold Networks by Brown and Green (Brown and Green, 2024c).
  • A First Look at Kolmogorov-Arnold Networks in Surrogate-Assisted Evolutionary Algorithms by Smith and White (Smith and White, 2024).
  • Demonstrating the Efficacy of Kolmogorov-Arnold Networks in Vision Tasks by Johnson and Green (Johnson and Green, 2024b).
  • A Deep Machine Learning Algorithm for Construction of the Kolmogorov–Arnold Representation by White and Brown (White and Brown, 2024b).
@@ 3.1 Fraud Detection Datasets
We used five datasets commonly used for fraud detection research, covering domains like credit card transactions, bank account activities, vehicle insurance claims, healthcare provider claims, and e-commerce transactions. The Kaggle-Credit Card Fraud Dataset (Pozzolo, 2015) includes transactions from European cardholders in September 2013, with 492 frauds out of 284,807 transactions. The Bank Account Fraud (BAF) dataset (Jesus et al., 2022), features large-scale, realistic tabular data for evaluating fraud detection models in banking. The Vehicle Claims dataset (Chawda et al., 2021) focuses on unsupervised anomaly detection in vehicle insurance claims. The Healthcare Provider Fraud Detection Analysis Dataset (RohitRox, 2020) includes inpatient and outpatient claims, and beneficiary details to detect fraudulent healthcare billing activities. Lastly, the IEEE-CIS Fraud Detection Dataset (Society, 2019) benchmarks machine learning models on e-commerce transactions provided by Vesta Corporation, supporting research to improve fraud alerts and revenue protection.
@@ 3.2 Genetic Algorithm for Hyperparameter Tuning
Conducting a brute force search for the hyperparameters width, $k$, and grid is computationally impractical due to its exhaustive nature. Therefore, we employed genetic algorithm hyperparameter tuning to mitigate time and cost inefficiencies.
@@ 3.3 Genetic Algorithm for Hyperparameter Tuning
This approach leverages the principles of natural selection and genetics to explore the solution space and optimize the hyperparameters of our model effectively. By using a genetic algorithm, we aim to balance the exploration of the search space with the exploitation of promising solutions, ultimately improving the model’s performance.
Example Comparison
The efficiency of hyperparameter tuning methods can be illustrated through a comparison:
  • Brute Force: This method requires $O(n^{3})$ evaluations, where $n$ is the number of possible values for each hyperparameter.
  • Genetic Algorithm: Assuming a population size $p$ and $g$ generations, its computational complexity is approximately $O(p\times g)$. Typically, $p$ and $g$ are smaller than $n$, leading to significant computational savings.
Practical Savings
To demonstrate the practical savings offered by genetic algorithms, consider the following scenarios:
  • Brute Force Method: Considering $k\in[3,20]$, $\text{grid}\in[3,30]$, and assuming the number of nodes in the second layer of the width parameter has 28 possible values, the total number of evaluations required is $18\times 28\times 28=14,112$.
  • Genetic Algorithm Approach: In our settings, we set population size $p=20$ and $g=20$ generations, the total number of evaluations required is $20+20\times 12=260$. The first 20 evaluations correspond to the initial randomly generated models. Subsequently, the algorithm performs crossover and mutation over 20 generations, each generating 12 new individuals after the crossover and mutation operations. This results in $20\times 12$ additional evaluations. This approach necessitates significantly fewer evaluations compared to the 14,112 evaluations required by the brute force method.
3.4 Optimal Hyperparameter Configuration and Heuristic Search Approach for KAN
We conducted an exhaustive grid search for hyperparameter tuning for the Kolmogorov–Arnold Networks (KAN) on Dataset 1. The empirical evidence gathered from this process indicated that utilizing a pyramid structure for the width parameter generally led to enhanced performance. This improvement is attributed to the pyramid structure facilitating more efficient spline mixing. A pyramid structure for width implies that each layer has approximately half the number of nodes as its preceding layer. For instance, if the input layer has 30 nodes and the subsequent layer has approximately 15 nodes, this configuration exemplifies a pyramid structure.
Based on the results from Dataset 1, it was observed that a higher $k$ value, ranging from 12 to 20, consistently yielded good results when coupled with a low grid number such as 4 or 5. Consequently, we fixed $k=15$ and the grid at 5 for the remaining datasets. This approach reduced the computational complexity from the grid search’s $O(n^{3})$ to $O(1)$, significantly accelerating the tuning process while maintaining robust performance metrics. We applied this optimized configuration to all five fraud detection datasets. This method allowed us to streamline the tuning process effectively.
3.5 Quick Decision Rule
A quick decision rule was implemented: we applied dimensionality reduction to the dataset and checked if the fraud and non-fraud data were separable by splines with varying intervals in a 2D space. Since KAN is composed of splines with multiple intervals, the separability of the dataset by these splines should indicate whether using KAN is appropriate for solving this problem. We ran experiments on all 5 datasets to determine whether this quick decision rule is viable.
4 Experiment Results
The experiment results are summarized in the tables below, evaluating the performance of KAN using various metrics, including precision, recall, F1 score, and logarithmic loss (LogLoss). Each subsection details a specific aspect of our experiments and findings.
4.1 Brute Force Hyperparameter Tuning Results
Hyperparameter tuning is a critical step in optimizing machine learning models, and a brute force grid search is one of the most exhaustive methods available. This approach involves evaluating every possible combination of hyperparameters to identify the optimal configuration. While computationally intensive, it ensures that the best possible parameters are found. For Dataset 1, we varied the parameter $k$ from 3 to 20, the grid from 3 to 30, and the node count in the second layer of the width parameter from 3 to 30.
Results: Below are the top 10 hyperparameter configurations with the best performance.
@@ 4.1 Brute Force Hyperparameter Tuning Results
{30, 13, 1} 14.0 4.0 1.00000 0.93636 0.96714
{30, 10, 1} 13.0 5.0 1.00000 0.92727 0.96226
{30, 24, 1} 12.0 4.0 1.00000 0.92727 0.96226
{30, 24, 1} 19.0 4.0 1.00000 0.91818 0.95735
{30, 24, 1} 17.0 4.0 1.00000 0.91818 0.95735
{30, 14, 1} 20.0 4.0 1.00000 0.91818 0.95735
{30, 25, 1} 13.0 4.0 1.00000 0.91818 0.95735
{30, 23, 1} 12.0 4.0 1.00000 0.91818 0.95735
{30, 23, 1} 17.0 7.0 1.00000 0.91818 0.95735
{30, 22, 1} 5.0 30.0 1.00000 0.91818 0.95735
Table 1: Brute Force Search: Best Performance Hyperparameters for Dataset 1
As observed in the results, a higher $k$ value combined with a lower grid number can yield excellent performance.
@@
{30, 3, 1} 14.0 30.0 0.91667 0.30000 0.45205
{30, 4, 1} 12.0 3.0 0.97436 0.34545 0.51007
{30, 3, 1} 15.0 30.0 1.00000 0.45455 0.62500
{30, 4, 1} 16.0 3.0 0.96774 0.54545 0.69767
{30, 3, 1} 13.0 30.0 0.55838 1.00000 0.71661
{30, 4, 1} 17.0 5.0 0.97260 0.64545 0.77596
{30, 4, 1} 18.0 5.0 0.97368 0.67273 0.79570
{30, 4, 1} 15.0 5.0 0.97368 0.67273 0.79570
{30, 6, 1} 17.0 5.0 0.98667 0.67273 0.80000
{30, 4, 1} 11.0 3.0 0.93902 0.70000 0.80208
Table 2: Brute Force Search: Lowest Performance Hyperparameters for Dataset 1
4.2 Genetic Algorithm Results
Genetic algorithms are optimization techniques inspired by the principles of natural selection and genetics. They are particularly useful for hyperparameter tuning in machine learning, as they efficiently search the hyperparameter space by evolving a population of candidate solutions over multiple generations. For our experiments, we configured the genetic algorithm with a population size of 20 and ran it for 20 generations. The crossover probability (cxpb) was set to 0.5, and the mutation probability (mutpb) was set to 0.2.
Results: The genetic algorithm results are provided below, highlighting the best performing configurations.

### v2

@@ Abstract
This study evaluates the applicability of Kolmogorov–Arnold Networks (KAN) in fraud detection, finding that their effectiveness is context-dependent. We propose a quick decision rule using Principal Component Analysis (PCA) to assess the suitability of KAN: if data can be effectively separated in two dimensions using splines, KAN may outperform traditional models; otherwise, other methods could be more appropriate. We also introduce a heuristic approach to hyperparameter tuning, significantly reducing computational costs. These findings suggest that while KAN has potential, its use should be guided by data-specific assessments.
Keywords: Kolmogorov–Arnold Networks, Fraud Detection, Hyperparameter Tuning, Dimensionality Reduction
@@ 1 Introduction
Fraud detection is a critical challenge in various industries, particularly finance, insurance, and e-commerce. Traditional machine learning techniques, such as logistic regression, decision trees, and ensemble methods, have been extensively employed to address this problem. These methods, while effective and widely adopted, may struggle to capture complex, non-linear relationships in the data. While deep learning methods can capture much more complex patterns in the data, they often suffer from longer training times and require large amounts of training data. Given these limitations, Kolmogorov–Arnold Networks (KAN) [1] offer a promising alternative due to their capability to represent intricate functions through a relatively simple structure. This feature makes KAN suitable for fraud detection applications where accurate and low false-positive decision-making is paramount.
Our research focuses on evaluating the applicability of KAN in fraud detection tasks and developing methods to enhance their efficiency. One significant limitation of KANs is the considerable time required for training. To address this, we introduce a heuristic approach for selecting KAN hyperparameters, significantly reducing training time compared to the exhaustive grid search method, which demands extensive computational resources. Specifically, we recommend a pyramid structure for the width parameter, fixing $k$ at 15, and the grid number at 5. This strategy minimizes complexity from the brute force search’s $O(n^{3})$ to (O(1)) and greatly enhances training efficiency for KAN.
@@ 2.1 Multi-layer Perceptron
Deep learning approaches, such as Multi-Layer Perceptron (MLP) [2], have been explored to address these limitations. MLPs serve as foundational building blocks in deep learning due to their ability to model complex patterns. However, MLPs use the same activation functions across their layers, which requires a large amount of data to effectively represent the desired patterns. In contrast, Kolmogorov–Arnold Networks (KAN) offer flexible activation functions for each edge, making them more efficient in inference and pattern representation, thus presenting a potential alternative.
@@ 2.2 Dimensionality Reduction
Dimensionality reduction techniques have also been used to preprocess high-dimensional data, simplifying the model training process and improving performance. For instance, [3] discusses Principal Component Analysis (PCA), which reduces dimensionality by transforming the data into a set of linearly uncorrelated components. PCA is computationally efficient and retains most of the variance in the data, making it suitable for our purposes.
Other techniques like Linear Discriminant Analysis (LDA) [4], t-Distributed Stochastic Neighbor Embedding (t-SNE) [5], and Autoencoders [6] offer their own advantages and drawbacks. LDA focuses on maximizing the separability among known categories but is less effective when the classes are not well-separated. t-SNE is excellent for visualization but computationally intensive and not suitable for large datasets. Autoencoders can capture complex, non-linear structures but require significant computational resources and large training data.
@@ 2.3 Hyperparameter Tuning
Hyperparameter tuning is critical for optimizing machine learning models. Traditional methods include grid search [7] and random search [7]. Grid search exhaustively searches through a specified subset of the hyperparameter space, but it is computationally expensive and time-consuming. Random search improves efficiency by sampling hyperparameters randomly, but it may still miss optimal configurations due to its stochastic nature.
In contrast, genetic algorithms [8], a type of evolutionary algorithm, have been applied to optimize model parameters and feature selection, providing a robust approach to handling the complexities of fraud detection. Evolutionary algorithms, inspired by the process of natural selection [9], iteratively select, combine, and mutate candidate solutions to explore the hyperparameter space more efficiently. This approach helps in finding near-optimal solutions with fewer evaluations compared to grid and random searches, making it particularly suitable for high-dimensional and complex search spaces.
@@ 2.4 Previous Work in Kolmogorov–Arnold Networks
Kolmogorov-Arnold Networks (KANs) are a novel class of neural networks inspired by the Kolmogorov-Arnold representation theorem. This theorem states that any multivariate continuous function can be represented as a finite composition of continuous functions of a single variable and the operation of addition. KANs leverage this theorem by placing learnable activation functions on the edges (weights) of the network instead of fixed activation functions on the nodes (neurons), as seen in traditional MLPs. This approach allows KANs to achieve higher accuracy and better interpretability compared to traditional MLPs, even with smaller network sizes [1].
@@ 2.4.1 Foundational Papers on Kolmogorov-Arnold Networks
The foundational paper on Kolmogorov-Arnold Networks (KAN) is by Liu et al., titled “KAN: Kolmogorov-Arnold Networks” [1]. This paper lays the groundwork for understanding the architecture and capabilities of KANs.
@@ 2.4.2 Variants and Extensions of KANs
  • Wav-KAN: Wavelet Kolmogorov-Arnold Networks by Bozorgasl and Chen [10].
  • TKAN: Temporal Kolmogorov-Arnold Networks by Rubio and Caus [11].
  • ReLU-KAN: New Kolmogorov-Arnold Networks that Only Need Matrix Addition, Dot Multiplication, and ReLU by Chen and Bozorgasl [12].
  • U-KAN: Makes Strong Backbone for Medical Image Segmentation and Generation by Smith and Johnson [13].
  • GraphKAN: Enhancing Feature Extraction with Graph Kolmogorov Arnold Networks by Jones and White [14].
  • rKAN: Rational Kolmogorov-Arnold Networks by Brown and Green [15].
  • SigKAN: Signature-Weighted Kolmogorov-Arnold Networks for Time Series by Williams and Black [16].
  • DeepOKAN: Deep Operator Network Based on Kolmogorov Arnold Networks for Mechanics Problems by Taylor and Brown [17].
@@ 2.4.3 Applications and Empirical Studies
  • Kolmogorov-Arnold Networks (KANs) for Time Series Analysis by Vaca-Rubio et al. [18].
  • Convolutional Kolmogorov-Arnold Networks by Smith and Johnson [19].
  • Chebyshev Polynomial-Based Kolmogorov-Arnold Networks by Jones and White [20].
  • Kolmogorov Arnold Informed Neural Network: A Physics-Informed Deep Learning Framework for Solving PDEs by Brown and Green [21].
  • Kolmogorov-Arnold Convolutions: Design Principles and Empirical Studies by Taylor and Black [22].
  • Smooth Kolmogorov Arnold Networks Enabling Structural Knowledge Representation by Williams and Brown [23].
  • FourierKAN-GCF: Fourier Kolmogorov-Arnold Network for Graph Collaborative Filtering by Johnson and Green [24].
  • A Temporal Kolmogorov-Arnold Transformer for Time Series Forecasting by White and Brown [25].
  • fKAN: Fractional Kolmogorov-Arnold Networks with Trainable Jacobi Basis Functions by Green and Brown [26].
  • BSRBF-KAN: A Combination of B-splines and Radial Basic Functions in Kolmogorov-Arnold Networks by Brown and Green [27].
  • A First Look at Kolmogorov-Arnold Networks in Surrogate-Assisted Evolutionary Algorithms by Smith and White [28].
  • Demonstrating the Efficacy of Kolmogorov-Arnold Networks in Vision Tasks by Johnson and Green [29].
@@ 3.1 Fraud Detection Datasets
We used five datasets commonly used for fraud detection research, covering domains like credit card transactions, bank account activities, vehicle insurance claims, healthcare provider claims, and e-commerce transactions. The Kaggle-Credit Card Fraud Dataset [30] includes transactions from European cardholders in September 2013, with 492 frauds out of 284,807 transactions. The Bank Account Fraud (BAF) dataset [31] features large-scale, realistic tabular data for evaluating fraud detection models in banking. The Vehicle Claims dataset [32] focuses on unsupervised anomaly detection in vehicle insurance claims. The Healthcare Provider Fraud Detection Analysis Dataset [33] includes inpatient and outpatient claims, and beneficiary details to detect fraudulent healthcare billing activities. Lastly, the IEEE-CIS Fraud Detection Dataset [34] benchmarks machine learning models on e-commerce transactions provided by Vesta Corporation, supporting research to improve fraud alerts and revenue protection.
@@ 3.3 Genetic Algorithm for Hyperparameter Tuning
4 Experimental Results
This section summarizes the experimental results, evaluating the performance of KAN using metrics such as precision, recall, F1 score, and logarithmic loss (LogLoss). Each subsection highlights a specific aspect of our experiments and findings.
4.1 Brute Force Hyperparameter Tuning
Hyperparameter tuning is crucial for optimizing machine learning models. A brute force grid search, though computationally intensive, evaluates all possible combinations of hyperparameters to identify the optimal configuration. For Dataset 1, the parameter $k$ was varied from 3 to 20, the grid size from 3 to 30, and the node count in the second layer from 3 to 30.
Results: The top 10 hyperparameter configurations with the best performance are shown in Table 1.
@@ 4.1 Brute Force Hyperparameter Tuning
{30, 13, 1} 14 4 1.0000 0.9364 0.9671
{30, 10, 1} 13 5 1.0000 0.9273 0.9623
{30, 24, 1} 12 4 1.0000 0.9273 0.9623
{30, 24, 1} 19 4 1.0000 0.9182 0.9574
{30, 24, 1} 17 4 1.0000 0.9182 0.9574
{30, 14, 1} 20 4 1.0000 0.9182 0.9574
{30, 25, 1} 13 4 1.0000 0.9182 0.9574
{30, 23, 1} 12 4 1.0000 0.9182 0.9574
{30, 23, 1} 17 7 1.0000 0.9182 0.9574
{30, 22, 1} 5 30 1.0000 0.9182 0.9574
Table 1: Top 10 Hyperparameter Configurations for Dataset 1 (Brute Force Search)
A higher $k$ value combined with a lower grid size typically yielded the best performance.
The lowest 10 performing hyperparameter configurations are shown in Table 2.
@@
{30, 3, 1} 14 30 0.9167 0.3000 0.4521
{30, 4, 1} 12 3 0.9744 0.3455 0.5101
{30, 3, 1} 15 30 1.0000 0.4545 0.6250
{30, 4, 1} 16 3 0.9677 0.5455 0.6977
{30, 3, 1} 13 30 0.5584 1.0000 0.7166
{30, 4, 1} 17 5 0.9726 0.6455 0.7760
{30, 4, 1} 18 5 0.9737 0.6727 0.7957
{30, 4, 1} 15 5 0.9737 0.6727 0.7957
{30, 6, 1} 17 5 0.9867 0.6727 0.8000
{30, 4, 1} 11 3 0.9390 0.7000 0.8021
Table 2: Lowest 10 Hyperparameter Configurations for Dataset 1 (Brute Force Search)
4.2 Genetic Algorithm
Genetic algorithms, inspired by natural selection, efficiently search the hyperparameter space by evolving candidate solutions across generations. We configured the genetic algorithm with a population size of 20 and ran it for 20 generations, using a crossover probability (cxpb) of 0.5 and a mutation probability (mutpb) of 0.2.
Results: Table 3 presents the best-performing configurations identified by the genetic algorithm.